# Laura Barros        Aiko Schmeißer

### v3

@@ Abstract
The supply of politicians affects the quality of democratic institutions. Yet little is known about the economic motivations that drive individuals into politics. This paper examines how experiencing a job loss affects individuals’ decisions to enter political life and explores its implications for political selection. Using highly granular administrative data linking individual records of political participation with comprehensive employer-employee data for all formal workers in Brazil, and leveraging mass layoffs for causal identification, we find that job loss significantly increases the likelihood of joining a political party and running for local office. Layoff-induced candidates are positively selected on various competence measures, suggesting that economic shocks improve the quality of political entrants. Further, we observe that the increase in candidacies is more pronounced among laid-off individuals with greater financial incentives from holding office and higher predicted income losses. Finally, using a regression discontinuity design, we find that eligibility for unemployment benefits increases the likelihood of becoming a party member and running for local office. These results are consistent with the reduction in private-sector opportunity costs and the increased time resources explaining the rise in political entry.
Keywords: Political selection, job loss, mass layoffs, unemployment benefits
@@ 1 Introduction
Politicians are perhaps the most important agents affecting the quality of democratic institutions. They aggregate citizens’ preferences in legislatures and influence which policies are enacted and how well they are implemented. Empirical work shows that political leaders strongly matter for the quality of policymaking (Pande, 2003; Jones and Olken, 2005; Besley et al., 2011; Clots-Figueras, 2012). But what motivates individuals to enter politics? Research on this question remains scarce and typically focuses on the role of institutional constraints – such as electoral rules and political reforms – for political selection (Besley et al., 2017; Arora, 2022; Avis et al., 2022). Little is known about the economic trade-offs that individuals face when deciding to join politics. Do political careers compete with careers in the private sector? In particular, does the private labor market draw talent away from political service? If individuals can substitute private and political jobs, changes in labor market opportunities may impact the candidates and representatives who select into political office. More broadly, negative economic shocks may have important consequences for the political system through the supply of politicians.
@@
In this paper, we study how experiencing a job loss affects individuals’ decision to enter politics and examine the consequences of these economic shocks for the quality of political candidates. Using rich administrative data from Brazil, we focus on party memberships and candidacies for local councils, the lowest-level positions in Brazilian politics. We rely on individual-level information on the universe of party members and local candidates between 2000 and 2020, which we match to employer-employee data on all formally employed workers between 2004 and 2018. The comprehensiveness of the data allows us to investigate effect heterogeneity along individual, local, and party characteristics and explore the consequences of eligibility for unemployment benefits after the job loss. In addition, we use various competence measures to document the implications of job loss for the selection of those who enter politics. Taken together, our results document novel insights into the mechanisms behind individual decisions to engage with politics following displacement from the private labor market, as well as the consequences of such economic shocks for the supply of politicians.
@@
To estimate the effects of job loss on the probability of entering politics, we exploit firm-level mass layoffs as plausibly exogenous variation for individual-level job loss. We compare individuals laid off in a mass layoff firm to a matched control group of similar workers from other firms who were not dismissed in the same year. The richness of our data allows us to exactly match treated and control workers on several characteristics, such as gender, education, age, wage, tenure, and firm size, and to control for fine-grained local- and industry-specific shocks that may jointly affect mass layoffs and political behavior. Importantly, accounting for changes in the local political environment ensures that our results capture individual supply responses and are not influenced by changes in the demand for politicians. We use a difference-in-differences (DiD) design to estimate dynamic treatment effects for up to three election cycles (12 years) after job loss. The underlying parallel-trends assumption is supported by the absence of differential trends in political outcomes between treated and control individuals before the layoff, and we also provide evidence that selection into (mass-)layoffs does not drive our results. [^1] Our empirical strategy closely follows recent papers studying the effects of mass layoffs in Brazil on crime (Britto et al., 2022b), domestic violence (Bhalotra et al., 2025), health (Amorim et al., 2025), and children’s education (Britto et al., 2022a).
@@
The positive effects of layoffs on political engagement are remarkably widespread. First, in robustness tests, we show that the effects do not differ strongly by the size of a mass layoff, mitigating concerns regarding external validity to regular layoffs. Second, we examine heterogeneity across a range of individual covariates (education, age, gender) and local characteristics (average income, unemployment rates, inequality, homicide rates, left vs. center-right presidential vote shares). We find significant positive effects on party memberships and candidacies in all the considered individual and local subsamples. Third, we show that the effects of job loss extend beyond the realms of party memberships and councilor candidacies, encompassing various facets of political participation. Specifically, we also find positive effects on the likelihood of running for mayor, donating to a political candidate, and working for a political campaign.
What explains the positive effects of job loss on political entry? Economic theories of political selection model the decision to run for office as a trade-off between the expected benefits from holding office and the costs of running (Black, 1972; Caselli and Morelli, 2004; Messner and Polborn, 2004). A layoff decreases the opportunity costs of entering politics by lowering the private-sector outside options and increasing the time available for political engagement. The size and granularity of our data enable us to examine heterogeneity in treatment effects and to test for the role of the opportunity cost channel in driving the increase in political entry after job loss. We start by exploiting individual variation in workers’ predicted income losses from the layoff and local variation in councilor salaries which depend on the size of Brazilian municipalities (Ferraz and Finan, 2011). The former captures the severity of the drop in private-sector opportunities predicted from individual observables and labor market characteristics, while the latter captures the financial attractiveness of a political career. We find larger layoff effects on candidacies among individuals who are predicted to face the greatest income loss due to a layoff. Consistently, the increase in candidacies is highest in municipalities with higher wage caps, with no corresponding effects for memberships. Together, these results show the importance of monetary incentives for individuals’ decision to enter politics following a job loss.
@@
An alternative explanation for the increase in political engagement following layoffs is that laid-off individuals may be mobilized through a change in party support. Individuals can be activated by parties promising to better represent their new economic policy preferences. Evidence suggests that more extreme parties, both on the left and the right, are often more successful in mobilizing voters who have experienced economic hardship (e.g., Algan et al., 2017; Autor et al., 2020). Moreover, the layoff may spark individuals into action as they seek to overturn the incumbent who is blamed for the change in their own economic circumstances (e.g., Tilley et al., 2018; Ahlquist et al., 2020). Another mobilizing force may arise from labor unions which often take part in mass layoff negotiations and have been shown to exert influence over the political behavior of employees (e.g., Kim and Margalit, 2017; Matzat and Schmeißer, 2023).
@@
A crucial question that follows is how layoffs affect the selection of politicians. Our results suggest that individuals’ decisions to engage with politics following layoffs are at least partly driven by economic rather than purely ideological motives. If economic incentives vary across individuals’ skill levels, economic shocks may also influence the average competence of politicians and, in turn, the functioning of the political system. To study the implications for the selection of political participants, we provide two pieces of evidence. First, we compare the characteristics of individuals who become party members and candidates due to the layoff (compliers) with those who would engage with politics regardless (always-takers) and those who would not engage with politics under any circumstances (never-takers). We find that compliers are positively selected in terms of education, previous wages, and AKM person wage effects, compared to always-takers and never-takers. For candidates, we also document that compliers have higher levels of education relative to the universe of candidates. Importantly, this does not come at the expense of political representation, as our results indicate that individuals who decide to run for office due to the layoff are younger and more likely to be female in comparison to always-takers and the population of candidates. Second, to study aggregate implications, we estimate the effects of municipal employment rates on the average education, gender, and age of local councilors. Exploiting local labor demand shocks from a Bartik (1991) instrument, we find that lower employment rates increase the average years of education not only among those individuals who run for office but also among those who are ultimately elected as councilors. Additionally, we observe that lower employment rates lead to a higher share of female and young councilors. Overall, these results suggest that economic downturns can lead to more positive political selection without compromising political representation.
Our paper contributes to two main strands of research. First, we speak to the literature that studies the determinants of political selection. [^3] See Dal Bó and Finan (2018) and Gulzar (2021) for recent reviews of this literature. In the tradition of the citizen-candidate framework (Osborne and Slivinski, 1996; Besley and Coate, 1997), this literature highlights the importance of endogenous entry into politics. Individuals’ decision to run for office is modeled as a cost-benefit calculation, where the expected benefits of holding office are weighed against the costs of running. On the benefit side, Ferraz and Finan (2011) and Gagliarducci and Nannicini (2013) document that increasing politicians’ salaries attracts higher-quality candidates and improves the performance of incumbent politicians. In addition to salaries, holding a public office may lead to other non-salary earnings while in office (whether legal or not) and enhance private sector opportunities after leaving office (Eggers and Hainmueller, 2009; Fisman et al., 2014). On the cost side, running for office may involve not only monetary, time, and opportunity costs, but also psychological and family burdens (Hall, 2019). Using detailed information on the social background of politicians in Sweden, Dal Bó et al. (2017) document that Swedish politicians are positively selected along various socio-demographic characteristics. In contrast, Britto et al. (2024) show that in Brazil, individuals with a criminal background are more likely to run for and win political office than the general population. Beyond individual determinants, political selection is also shaped by institutional factors, such as electoral and fiscal rules (Arora, 2022; Gamalerio and Trombetta, 2025), gender quotas (Besley et al., 2017), campaign spending limits (Avis et al., 2022), and political training programs (Dal Bó et al., 2024).
Our contribution to this line of research is to provide evidence on how economic trade-offs motivate individuals to enter politics. Specifically, we document how an adverse shock in private-sector opportunities increases individuals’ likelihood of starting a political career. The effects increase with several measures of individual competence, suggesting that layoffs can improve candidate quality. In addition, we show that unemployment benefits, by changing individuals’ time constraints after layoffs, impact political selection, resulting in an increased entry of better-educated individuals into politics. To the best of our knowledge, we are the first to study the impact of layoffs on political entry using large-scale individual-level data, which allows us to tackle the rare-event problem that researchers typically face in the study of political selection.
@@
Second, we add to the literature on the political consequences of economic shocks. A growing number of papers document that voting results respond to macro shocks, such as import competition (Autor et al., 2020; Dippel et al., 2022), economic crisis (Barros and Santos Silva, 2025), automation (Anelli et al., 2019), and austerity reforms (Fetzer, 2019). Similarly, at the individual level, many studies document that changes in economic conditions affect individuals’ policy preferences, including support for redistribution and welfare provision (e.g. Margalit, 2013; Martén, 2019; Ahlquist et al., 2020; Ballard-Rosa et al., 2021). In terms of political participation, the literature has focused mostly on turnout and has found mixed results, ranging from turnout increasing with local unemployment rates (Charles and Stephens, 2013; Burden and Wichowsky, 2014), decreasing after individual job loss (Emmenegger et al., 2017; Östermann and Lindgren, 2023), or not being affected at all by income changes (Jungkunz and Marx, 2022; Geys and Sørensen, 2025).
@@
We contribute to this strand of literature by revealing that economic shocks not only affect the demand of voters, but also impact the supply side of politics. In particular, we document how layoffs increase individuals’ propensity to enter the political system, either as members of a political party or as candidates in local elections. In addition, our mass-layoff identification strategy, together with the size and granularity of our data, allows us to exploit exogenous changes in personal economic conditions and study the mechanisms behind individuals’ decisions to engage with politics after a layoff.
@@ 2 Institutional and Theoretical Background
In Brazil, a significant portion of the population is engaged in politics. As mandated by the 1988 Brazilian constitution, voting is compulsory for all literate citizens aged 18 to 69.444The average turnout rate in Federal Elections in the last 20 years was about 80% (International Institute for Democracy and Electoral Assistance, 2024). Apart from voting, more than 16 million Brazilians were registered as members of a political party in 2020. This corresponds to about 11% of the electorate, among the highest shares observed across democratic countries. In addition, between 2000 and 2020, about 1.7 million individuals ran for office in local elections in Brazil. In this section, we describe the institutional environment in which individuals decide about their involvement as party members or local candidates and present a theoretical framework that illustrates the role of opportunity costs in this decision.
@@ 2.1 Party members
Brazil has a multi-party political system, encompassing 33 registered parties as of 2024. The effective functioning of political parties relies heavily on an engaged membership base. According to the Brazilian electoral law, a party must have a minimum number of members equal to 0.5% of the votes cast in the previous national election. Challenging the notion that political affiliation in Brazil is solely driven by clientelistic motives, Ribeiro and Do Amaral (2019) show that most political parties are well-organized and have a strong base of activists. Members carry out the day-to-day organizational work, participate in electoral campaigns, raise funding, attend party meetings, help maintain ties with the electorate outside of election years, and serve as the pool from which parties recruit their candidates and officials.
@@
Membership recruitment in Brazil is predominantly a local phenomenon, fueled by internal competition at parties’ municipal conventions. Local elites, aiming to secure the party’s nomination, attract supporters by enlisting them in the party to build a loyal voting base at the convention and signal their potential as candidates (Sells, 2020; Frey, 2022). Supporters who become members of a winning political party are often rewarded with employment opportunities within the government bureaucracy (Brollo et al., 2017; Barbosa and Ferreira, 2023). As suggestive evidence of such benefits in our setting, Appendix Figure A.1 shows that laid-off individuals are more likely to start a public-sector job if they were member of a political party before the layoff, with no comparable increase observed for private-sector employment. The strategic importance of party members in local politics is further reflected in the cyclical pattern of new affiliations: as shown in Appendix Figure A.2, parties significantly expand their ranks in the year before local elections.

### v4

@@ Abstract
Adverse economic shocks are known to reshape voter behavior – the demand side of politics. Much less is known about their consequences for the supply side: how such shocks affect who becomes a politician. This paper examines how job losses influence individuals’ decisions to enter politics and the implications for political selection. Using administrative data linking political participation records to matched employer-employee data covering all formal workers in Brazil, and exploiting mass layoffs for causal identification, we find that job loss significantly increases the likelihood of joining a political party and running for local office. Layoff-induced candidates are positively selected on various competence measures, indicating that economic shocks can improve the quality of political entrants. The increase in candidacies is strongest among laid-off individuals with greater financial incentives from holding office and higher predicted income losses. A regression discontinuity design further shows that eligibility for unemployment benefits increases political entry. These results are consistent with a reduction in individuals’ opportunity costs – both in terms of reduced private-sector income and increased time resources – facilitating greater political engagement.
Keywords: Political selection, economic shocks, job loss, mass layoffs
@@ 1 Introduction
A large empirical literature documents that adverse economic conditions can undermine democratic stability: rising unemployment and declining incomes reduce trust in institutions, exacerbate political polarization, and increase support for populist and extremist parties (e.g., Funke2016; Algan2017a; Autor; Guiso2023). This research has primarily focused on the demand side of democratic erosion – that is, how economic hardship reshapes the political preferences and electoral behaviors of voters. Yet an equally important, but less explored, dimension concerns the supply side of politics: How do economic shocks shape individuals’ incentives and opportunities to become a politician?
Politicians are perhaps the most important agents affecting the quality of democratic institutions. They aggregate citizens’ preferences in legislatures and influence which policies are enacted and how well they are implemented (e.g., Pande2003; Jones2005; Besley2011; ClotsFigueras2012). Understanding how economic conditions influence who enters politics is therefore crucial. Individuals considering a political career face opportunity costs that depend on conditions in the private labor market. When private sector opportunities are abundant, talented individuals may be drawn away from political service. Conversely, when these opportunities contract, political careers may become more attractive, potentially increasing both the number and the quality of those who choose to seek office. Thus, in contrast to their role for the demand side of politics, adverse economic shocks may have positive consequences for the political system by improving the supply of politicians.
@@
In this paper, we shed new light on the political consequences of economic shocks by studying how experiencing a job loss affects individuals’ decision to enter politics. Using rich administrative data from Brazil, we focus on party memberships and candidacies for local councils, the lowest-level positions in Brazilian politics. We rely on individual-level information on the universe of party members and local candidates between 2000 and 2020, which we match to employer-employee data on all formally employed workers between 2004 and 2018. Various individual competence measures allow us to document the implications of job loss for the selection of those who enter politics. Moreover, we investigate effect heterogeneity along individual, local, and party characteristics and explore the consequences of eligibility for unemployment benefits after the job loss, providing novel insights into the mechanisms behind individual decisions to engage with politics following the displacement.
@@
To estimate the effects of job loss on the probability of entering politics, we exploit firm-level mass layoffs as plausibly exogenous variation for individual-level job loss. We compare individuals laid off in a mass layoff firm to a matched control group of similar workers from other firms who were not dismissed in the same year. The richness of our data allows us to exactly match treated and control workers on several characteristics, such as gender, education, age, wage, tenure, and firm size, and to control for fine-grained local- and industry-specific shocks that may jointly affect mass layoffs and political behavior. Importantly, accounting for changes in the local political environment ensures that our results capture individual supply responses and are not influenced by changes in the local demand for politicians. We use a difference-in-differences (DiD) design to estimate dynamic treatment effects for up to three election cycles (12 years) after job loss. The underlying parallel-trends assumption is supported by the absence of differential trends in political outcomes between treated and control individuals before the layoff, and we also provide evidence that selection into (mass-)layoffs does not drive our results. [^1] Our empirical strategy closely follows recent papers studying the effects of mass layoffs in Brazil on crime (Britto2022), domestic violence (Bhalotra2021), health (Amorim2023), and children’s education (Britto2022a).
@@
The positive effects of layoffs on political engagement are remarkably widespread. First, in robustness tests, we show that the effects do not differ strongly by the size of a mass layoff, mitigating concerns regarding external validity to regular layoffs. Second, we examine heterogeneity across a range of individual covariates (education, age, gender) and local characteristics (average income, unemployment rates, inequality, homicide rates, left vs. center-right presidential vote shares). We find significant positive effects on party memberships and candidacies in all the considered individual and local subsamples. Third, we show that the effects of job loss extend beyond the realms of party memberships and councilor candidacies, encompassing various facets of engagement in politics. Specifically, we also find positive effects on the likelihood of running for mayor, donating to a political candidate, and working for a political campaign.
What explains the positive effects of job loss on political entry? Economic theories of political selection model the decision to run for office as a trade-off between the expected benefits from holding office and the costs of running (Black1972; Caselli2004; Messner2004). A layoff decreases the opportunity costs of entering politics by lowering the private-sector outside options and increasing the time available for political engagement. The size and granularity of our data enable us to examine heterogeneity in treatment effects and to test for the role of the opportunity cost channel in driving the increase in political entry after job loss. We start by exploiting individual variation in workers’ predicted income losses from the layoff and local variation in councilor salaries which depend on the size of Brazilian municipalities (Ferraz2011). The former captures the severity of the drop in private-sector opportunities predicted from individual observables and labor market characteristics, while the latter captures the financial attractiveness of a political career. We find larger layoff effects on candidacies among individuals who are predicted to face the greatest income loss due to a layoff. Consistently, the increase in candidacies is highest in municipalities with higher wage caps, with no corresponding effects for memberships. Together, these results show the importance of monetary incentives for individuals’ decision to enter politics following a job loss.
@@
An alternative explanation for the increase in political engagement following layoffs is that laid-off individuals may be mobilized through a change in party support. Individuals can be activated by parties promising to better represent their new economic policy preferences. Evidence suggests that more extreme parties, both on the left and the right, are often more successful in mobilizing voters who have experienced economic hardship (e.g., Algan2017a; Autor). Moreover, the layoff may spark individuals into action as they seek to overturn the incumbent who is blamed for the change in their own economic circumstances (e.g., Tilley2018; Ahlquist2020). Another mobilizing force may arise from labor unions which often take part in mass layoff negotiations and have been shown to exert influence over the political behavior of employees (e.g., kim2017; matzat2023).
@@
A crucial question that follows is how layoffs affect the selection of politicians. Our results suggest that individuals’ decisions to engage with politics following layoffs are at least partly driven by economic rather than purely ideological motives. If economic incentives vary across individuals’ skill levels, economic shocks may also influence the average competence of politicians and, in turn, the functioning of the political system. To study the implications for the selection of political participants, we provide two pieces of evidence. First, we compare the characteristics of individuals who become party members and candidates due to the layoff (compliers) with those who would engage with politics regardless (always-takers) and those who would not engage with politics under any circumstances (never-takers). We find that compliers are positively selected in terms of education, previous wages, and AKM person wage effects, compared to always-takers and never-takers. For candidates, we also document that compliers have higher levels of education relative to the universe of candidates. Importantly, this does not come at the expense of political representation, as our results indicate that individuals who decide to run for office due to the layoff are younger and more likely to be female in comparison to always-takers and the population of candidates. Second, to study aggregate implications, we estimate the effects of municipal employment rates on the average education, gender, and age of local councilors. Exploiting local labor demand shocks from a bartik1991 instrument, we find that lower employment rates increase the average years of education not only among those individuals who run for office but also among those who are ultimately elected as councilors. Additionally, we observe that lower employment rates lead to a higher share of female and young councilors. Overall, these results suggest that economic downturns can lead to more positive political selection without compromising political representation.
@@
Our paper contributes to two main strands of research. First, we speak to the literature on the political consequences of economic shocks. A large body of work shows that voting outcomes are influenced by macro shocks, such as import competition (colantone2018; Autor; dippel2022), economic crisis (deBromhead2013; Funke2016; Barros2019), automation (Frey2018; anelli2019; Im2019), and austerity policies (Fetzer2019; GalofreVila2021; DuqueGabriel2025). [^3] For a comprehensive review of the literature linking economic shocks to the rise of populism, see Guriev2022. Similarly, at the individual level, many studies document that changes in economic conditions affect citizens’ policy preferences, including support for redistribution and welfare provision (margalit2013; marten2019; Ahlquist2020; ballardrosa2021). In terms of political participation, the literature has focused mostly on voter turnout and has found mixed results, ranging from turnout increasing with local unemployment rates (charles2013; burden2014), decreasing after individual job loss (emmenegger2017; oestermann2023), or not being affected at all by income changes (jungkunz2022; geys2024).
@@
We add to this strand of literature by revealing that economic shocks affect not only the demands of voters, but also impact the supply side of politics. Specifically, we document that layoffs increase individuals’ propensity to enter the political system, both as members of a political party and as candidates in local elections. The effects increase with several measures of individual competence, suggesting that negative labor market shocks can improve the quality of political entrants. To the best of our knowledge, we are the first to study the causal impact of changes in individual economic conditions on political participation using administrative data and an identification strategy that exploits individuals’ exposure to exogenous layoff shocks.
Second, we contribute to the literature on the determinants of political selection. [^4] See besley2005, DalBo2018 and Gulzar2021 for reviews of this literature. In the tradition of the citizen-candidate framework (osborne1996; besley1997), this literature highlights the endogenous nature of entry into politics. Descriptive studies document how politicians differ from the general population: in Sweden, office holders are positively selected along various competence measures (DalBo2017), while in Brazil, politicians are more likely to have a criminal background (britto2024). Causal analyses of political selection focus mostly on the role of institutional determinants, such as electoral and fiscal rules (arora2022; beath2016; gamalerio2025), gender quotas (besley2017), campaign spending limits (avis2022), and political training programs (dal2024). Most closely related to our paper, Ferraz2011 and gagliarducci2013 find that higher politician salaries attract more educated candidates for local political offices in Brazil and Italy, respectively. Both studies demonstrate how financial incentives influence who enters politics by altering the returns to holding office.
We extend these studies by examining the reverse mechanism: how reductions in private-sector opportunities – and thus reductions in the opportunity cost of holding office – affect political entry. In doing so, we provide novel evidence on the economic trade-offs that motivate individuals to engage in politics. Leveraging rich administrative data, we shed light on the channels through which layoffs increase political entry: they lower individuals’ opportunity costs in terms of foregone income and time constraints. We further show that layoffs improve the selection of political candidates along several competence measures, and that access to unemployment benefits amplifies the entry of better-educated individuals into politics. Finally, the large-scale individual-level data allow us to tackle the rare-event problem that typically constrains empirical work on political selection.
@@ 2 Institutional and Theoretical Background
In Brazil, a significant portion of the population is engaged in politics. As mandated by the 1988 Brazilian constitution, voting is compulsory for all literate citizens aged 18 to 69.555The average turnout rate in Federal Elections in the last 20 years was about 80% (IDEA2024). Apart from voting, more than 16 million Brazilians were registered as members of a political party in 2020. This corresponds to about 11% of the electorate, among the highest shares observed across democratic countries. In addition, between 2000 and 2020, about 1.7 million individuals ran for office in local elections in Brazil. In this section, we describe the institutional environment in which individuals decide about their involvement as party members or local candidates and present a theoretical framework that illustrates the role of opportunity costs in this decision.
@@ 2.1 Party members
Brazil has a multi-party political system, encompassing 33 registered parties as of 2024. The effective functioning of political parties relies heavily on an engaged membership base. According to the Brazilian electoral law, a party must have a minimum number of members equal to 0.5% of the votes cast in the previous national election. Challenging the notion that political affiliation in Brazil is solely driven by clientelistic motives, ribeiro2019 show that most political parties are well-organized and have a strong base of activists. Members carry out the day-to-day organizational work, participate in electoral campaigns, raise funding, attend party meetings, help maintain ties with the electorate outside of election years, and serve as the pool from which parties recruit their candidates and officials.
@@
Membership recruitment in Brazil is predominantly a local phenomenon, fueled by internal competition at parties’ municipal conventions. Local elites, aiming to secure the party’s nomination, attract supporters by enlisting them in the party to build a loyal voting base at the convention and signal their potential as candidates (sells2020; frey2022). Supporters who become members of a winning political party are often rewarded with employment opportunities within the government bureaucracy (brollo2017; barbosa2023). As suggestive evidence of such benefits in our setting, Appendix Figure A.1 shows that laid-off individuals are more likely to start a public-sector job if they were members of a political party before the layoff, with no comparable increase observed for private-sector employment. The strategic importance of party members in local politics is further reflected in the cyclical pattern of new affiliations: as shown in Appendix Figure A.2, parties significantly expand their ranks in the year before local elections.